# Berlin, Germany

### v2

@@ 3.2.2 Mahalanobis Distance and Fitting Data
Let $B$ denote a set of samples $b^{1},\cdots,b^{l}$ from a probability distribution in $\mathcal{M}$ with mean $\mu$. The matrix $\frac{1}{l-1}\sum_{j=1}^{l}\log_{\mu}b^{j}(\log_{\mu}b^{j})^{T}$ (the superscript $T$ represents the transpose) in exponential chart at $\mu$ defines an endomorphism $\Sigma$ on $T_{\mu}\mathcal{M}$, which is independent of the coordinates (cf. [21]). Thus, $\Sigma$ extends the the notion of ordinary Euclidean (sample) covariance matrix to general Riemannian manifolds. Its inverse $S:=\Sigma^{-1}$ is the precision operator. The squared Mahalanobis distance between $B$ and an arbitrary point $x\in M$ denoted by $d_{Mah}$ reads
@@
Now, let $j=1,\cdots,l$ and consider a set $Y$ of discrete trajectories (not necessarily with equal sample sizes) $y^{j}$ in $M$ and $B$ be the set of corresponding control points $b^{j}=(b_{1}^{j},\cdots,b_{n}^{j})\in\mathcal{M}$ of the best-fitting polynomials for the trajectories, gained by minimizing $H$. Now, we may write
$F_{\mu,\lambda S}(b)$ $=H(b)+\lambda d_{Mah}^{2}(B,b).$
Thus, $F_{\mu,\lambda S}$ averages the least-squares regression using $H$ given in the previous subsection and the squared Mahalanobis distance to $B$.
@@ 3.2.3 Parameter Optimization and Forecasting
Now, fix $\lambda$ and initial samples $y_{1},\cdots,y_{i-1}$ of a discrete trajectory not in the prior. Starting with $i=i_{0}$, we aim to predict $y_{i}$. Our initial prediction reads $\hat{y}_{i}:=p(t_{i};b_{i}^{\ast})$, where $b_{i}^{\ast}$ denotes the minimizer of $F_{\mu,\lambda S}$. A common issue is that, in general, for coarse sampling as well as for long-term forecasting ($t_{i}-t_{i-1}$ much larger than the preceding time step), the error $d(\hat{y}_{i},y_{i})$ can increase excessively. To account for such effects and fine tuning, we slightly extend our ridge regression by adding a simple and computationally fast averaging step as follows. We replace $\hat{y}_{i}$ with $\exp_{y_{i-1}}(\alpha\log{y_{i-1}}\hat{y}_{i})$. Note that for $\alpha=1$, $\hat{y}_{i}$ remains unchanged. Iteration over $i$ yields the initial prediction for the whole trajectory. The initial and final values of $i$ are application dependent.
@@
Our next task is to find optimal values for the parameter $\lambda$ and $\alpha$. To this end, one can use the well-known methods in the Euclidean case such as cross-validation. Thereby, denoting the validation dataset by $\tilde{Y}$, optimal parameter values $(\lambda^{\ast},\alpha^{\ast})$ can be gained iteratively via minimization of the residual
$\sum_{\tilde{y}=(\tilde{y}_{1},\cdots,\tilde{y}_{m})\in\tilde{Y}}\sum_{i=i_{0}}^{m}d^{2}(\hat{y}_{i},\tilde{y}_{i}).$
@@
Obviously, the choice of the prior and its splitting in $\tilde{Y}$ and $Y$ depends on the application and particularly, properties like seasonality and relevance of historical order of the trajectories. Moreover, using a larger prior requires more computation time and memory. However, it does not necessarily result in better prediction accuracy on the test dataset. On the other hand, it is unlikely to achieve satisfactory prediction accuracy with smaller $Y$ and $\tilde{Y}$.
4 Application: Hurricanes Forecast
Tropical cyclones, also known as hurricanes or typhoons, are among the most powerful natural phenomena with enormous environmental, economic, and human impact. The most common indicator of a hurricane’s intensity is its maximum sustained wind speed, which is used to categorize the storm on the Saffir-Simpson hurricane wind scale. For example, wind speeds $\geq 137$ knots correspond to category 5. High track variability and out-most complexity of hurricanes has led to a large number of works to classify, rationalize and predict them. We remark that many approaches are not intrinsic and use linear approximations. We refer to the overview [10], the summary of recent progress [7], and [11].
In [16], we represented hurricane tracks using Bézier splines (compositions of Bézier polynomials), computed average trajectories, performed tangent principal component analysis, and presented an intensity classification using support vector machines. In the following, we apply our proposed ridge regression approach, based on Bézier polynomial representation, to hurricane forecasting.
4.1 Dataset
We verify the effectiveness of the proposed framework by applying it to the Atlantic hurricane data from the HURDAT 2 database provided by the U.S. National Oceanic and Atmospheric Administration, publicly available on https://www.nhc.noaa.gov/data/. The data comprise measurements of latitude, longitude, and wind speed on a 6 hours base. The sample size (number of points that make up a track) varies from 13 to 96. Minimum of observed wind speeds is $15$ knots. Figure 1 illustrates this dataset depicting the 2021 tracks together with their intensities.
4.2 Experiments and Discussion
In our experiments, following the verification rules of the U.S. National Hurricane Center (NHC), we computed the average of the MAE (mean absolute error; using spherical distance for the tracks) of the 2021 Atlantic forecasts for comparison. To this end, we represented the tracks and intensities as discrete trajectories in $S^{2}$ and $\mathbb{R}$, respectively. We recall that the Riemannian exponential and logarithmic maps for $S^{2}$ read as follows.
@@ 4.2 Experiments and Discussion
where $x,y\in S^{2}$ with $dist(x,y)<\pi$ and $v$ orthogonal to $x$ with length $\varphi$ ($\exp_{x}0=x$). We initiated each forecast using the first sample (initial value of $i$ was $2$) and choose $n=6$, resulting in average $R^{2}$ values $>0.85$ for both fitted tracks and intensities. Figure 2 shows two example tracks with their representation in our polynomial regression model.
Figure 2: Two exemplary tracks (blue) with their representation (red) in our polynomial regression model.
@@
We employed two experiments, Exp1 and Exp2. Due to the deviations of annual trends and relevance of the chronological order, in Exp1, we used the 2020 ones (31 trajectories) as prior and the last 21 for validation to forecast all the 2021 ones (21 trajectories), and in Exp2, the first 16 trajectories from 2021 as prior with its last 5 trajectories to forecast the last 5 trajectories. It turned out that while the intensities are highly correlated, the covariance matrix of the tracks is almost singular. We solved the latter issue employing diagonal loading (simple standard approach to numerically treat matrices having determinant close to zero, i.e., $\approx$ machine precision; cf. [5]). Higher values of $n$ increased the forecast error of the tracks (overfitting due to small position correlations). For parameter optimization we only employed a simple rough grid search to decrease the computation time (on average few seconds for each forecast). Variations in the settings with significantly larger or smaller validation sets increased the errors, and as expected, validation sets chronologically closer to the test sets resulted in forecasts with slightly higher accuracies.
@@
Obviously, one can model the tracks as images of the Euclidean best-fitting polynomials in geographic coordinates (in the latitude-longitude plane) under the standard parametrization of the sphere. Our computations show that this increases errors, demonstrating the advantage of our intrinsic approach, which incorporates the geometry of the data. Figure 3 shows an example with a particularly large track error, which is mainly caused by the loop. The increase in error due to the presence of a loop, in addition to meteorological anomalies, is an issue known in other approaches (cf.  [2]).
@@
Figure 3: 12 h forecasts (red) via ridge regression for the category 4 and long-lasting 2021 hurricane Sam (blue).
Table 1 summarizes the average 12 h forecast errors of the standard NHC model OCD5 (cf. [9]) and our two primary experiments regarding the Atlantic basin for the 2021. Of course, we did not expect results very close to those of NHC, since our approach is generic and does not include all the relevant physical parameters, while NHC uses several more elaborate physical models. However, our approach provides a flexible framework that can complement existing physical models, especially in scenarios where physical data is sparse, computationally intensive, or when exploring shape and morphological features. Moreover, it lays a foundation for future enhancements, such as incorporating additional physical parameters that could further improve predictive performance.
Table 1: Average 12 h forecast errors of NHC/OCD5 model and proposed approach in the Atlantic basin for the 2021 (Exp1: all, Exp2: last 5) cyclones.
Experiment/Method Intensities (kt) Tracks (mi)
Exp1/NHC 6.4 48.2
Exp1/Proposed 6.9 90.7
Exp2/NHC 7.5 67.4
Exp2/Proposed 6.5 97.2
@@
For the code implementing our approach, which in particular includes Riemannian optimization for the computation of geodesic paths and Bézier polynomials, we used the publicly available Python package morphomatics v4.0 (cf. [3]).
5 Conclusion
@@
In this work, we presented an intrinsic natural extension of polynomial ridge regression, respectively, Tikhonov regularization, from Euclidean spaces to arbitrary Riemannian manifolds. This allows for the prediction of manifold-value time series using the Mahalanobis distance to the training data as a prior. We also provided formulae for computing the gradient of the loss function in the approach. In addition, we provided a discussion of our numerical experiments on the application of the approach to hurricane forecasting and intensity.

### v3

@@ 2.2.2 Mahalanobis Distance and Fitting Data
Let $B$ denote a set of samples $b^{1},\cdots,b^{l}$ in $\mathcal{M}$ with mean $\mu$. The matrix $\frac{1}{l-1}\sum_{j=1}^{l}\mathrm{Log}_{\mu}b^{j}(\mathrm{Log}_{\mu}b^{j})^{T}$ (the superscript $T$ represents the transpose) in the exponential chart at $\mu$ defines an endomorphism $\Sigma$ on $T_{\mu}\mathcal{M}$, which is independent of the coordinates (cf. [pennec2006intro]). Thus, $\Sigma$ extends the notion of ordinary Euclidean (sample) covariance matrix to general Riemannian manifolds. Its inverse $S:=\Sigma^{-1}$ is the precision operator. The squared Mahalanobis distance between $B$ and an arbitrary point $x\in M$ denoted by $d_{Mah}$ reads
@@
Thus, the Mahalanobis distance accounts for data distribution. Now, consider a set $Y$ of discrete trajectories (not necessarily with equal sample sizes) in $M$ and let $B\subset\mathcal{M}$ be the set of corresponding control points of the best-fitting polynomials for the trajectories in $Y$, obtained by minimizing $H$. We may write
$F_{\lambda}(b)$ $=H(b)+\lambda d_{Mah}^{2}(B,b).$
Thus, $F_{\lambda}$ averages the least-squares regression using $H$ given in the previous subsection and the squared Mahalanobis distance to $B$.
@@ 2.2.3 Parameter Optimization and Forecasting
Now, fix $\lambda$ and initial samples $y_{1},\cdots,y_{m}$ of a discrete trajectory. We aim to predict $y_{m+1}$. Our prediction reads $\hat{y}_{m+1}:=p(t_{m+1};b^{\ast})$, where $b^{\ast}$ denotes the minimizer of $F_{\lambda}$. A common core issue in forecasting is that, in general, for coarse sampling as well as for long-term forecasting ($t_{i}-t_{i-1}$ much larger than the preceding time step), the error $d(\hat{y}_{i},y_{i})$, due to noise and error propagation can increase excessively. To account for such effects and fine tuning, we slightly extend our ridge regression by adding a simple and computationally fast averaging step as follows. To this end, we consider an extended predictor $P_{\alpha}$
$P_{\alpha}(t;b,x)=\exp_{x}(\alpha\log_{x}(p(t;b)))$
with $\alpha\in[0,1]$ and replace $\hat{y}_{m+1}$ by $P_{\alpha}(t_{m+1};b^{*},y_{m})$. Note that for $\alpha=1$, $\hat{y}_{m+1}$ remains unchanged. Iteration over $m$ yields the prediction for the whole trajectory.
@@
Now, our task is to find optimal values for the parameter $\lambda$ and $\alpha$. To this end, one can use the well-known methods in the Euclidean case such as cross-validation. Thereby, denoting the validation dataset by $Y_{val}$, optimal parameter values $(\lambda^{\ast},\alpha^{\ast})$ can be gained iteratively via minimization of the residual
$\sum_{y=(y_{1},\cdots,y_{m})\in Y_{val}}\sum_{i=1}^{m}d^{2}(\hat{y}_{i},y_{i}).$
@@
Obviously, the choice of the prior and its splitting in $Y_{val}$ and $Y$ depends on the application and, particularly, on properties such as seasonality and relevance of the historical order of the trajectories. Moreover, using a larger prior requires more computation time and memory. However, it does not necessarily result in better prediction accuracy on the test dataset. However, it is unlikely to achieve satisfactory prediction accuracy with smaller $Y$ and $Y_{val}$. Note that the ridge parameter $\lambda$ controls the tradeoff between complexity and fit by penalizing solutions that are too far from the mean $\mu$. For details on the complete forecasting pipeline, see Appendix 5.3.
3 Numerical Experiments
We demonstrate the effectiveness of our approach through two validations: (1) controlled synthetic experiments on spherical trajectories where ridge regularization reduces prediction error compared to unregularized regression, and (2) real-world hurricane forecasting where our method achieves competitive performance using only historical data.
In our experiments, $M$ was the 2-sphere embedded in $R^{3}$. We recall that the Riemannian exponential and logarithmic maps for $S^{2}$ read
@@ 3 Numerical Experiments
where $x,y\in S^{2}$ with $dist(x,y)<\pi$, $v\in T_{x}S^{2}$ with $\|v\|=\varphi$ and $\exp_{x}0=x$. The computed $3n\times 3n$ covariance matrix in ambient coordinates is singular by construction, as all tangent vectors $\mathrm{Log}_{\mu}b$ lie in the $2n$-dimensional space $T_{\mu}(S^{2})^{n}$. We employed Principal Component Analysis for dimensionality reduction, retaining only principal components with positive eigenvalues, effectively capturing the data’s intrinsic dimensionality. For comparison, simpler regularized estimation methods, specifically Ad Hoc Diagonal Loading [diagload2005performance] and the Ledoit-Wolf technique [LEDOIT2004365], were also applied. In all cases, these alternative regularization methods yielded results statistically equivalent to those obtained using the PCA-derived effective covariance matrix, confirming the robustness of the results against the specific choice of covariance estimator for this dataset.
3.1 Synthetic Validation
A natural question is whether ridge regression on manifolds provides advantages compared to the ordinary least squares (OLS), $\lambda=0$, approach as known in the Euclidean case. Towards answering this question, we employed the following experiments with spherical data. We generated three template curves (mean trajectories): a geodesic, a cubic Bézier curve and the image of a perturbed sinusoidal curve in latitude-longitude coordinates, each consisting of 305030--50 points, and conducted 3 experiments ExpGeo, ExpCubic and ExpSin respectively. In each experiment, we generated 30 noisy trajectories by adding spatially-correlated random Gaussian noise in the tangent space at each point, with noise standard deviation $0.05$, followed by projection via the exponential map. This construction ensured between-trajectory correlation while maintaining realistic variability. Thus, in each case the resulting dataset comprised 30 trajectories, which were randomly split into 20 for training, 5 for validation and 5 for testing. We choose degree-5 Bézier polynomials yielding average $R^{2}$ value of $95\%$ to ensure high goodness of model-fit. For prediction, we set $\alpha=1$ and employed an iterative forecasting scheme where at each time step, we used the last $6$ observations to fit the model and predict the next position.Using the optimal value $\lambda^{*}$ obtained by cross-validation yielded a minimum average MAE (mean absolute error) in all experiments smaller than the one from OLS ($\lambda=0$), representing a $21.4\%$ average improvement. This demonstrates that ridge regularization on manifolds can provide substantial benefits when the data exhibit correlation structure, analogous to the Euclidean setting. Exemplary trajectories and their forecasts are depicted in Figure 2, and summary results are presented in Table 1.
Figure 2: Two Representative forecasts from synthetic experiments, illustrating the prediction of the ridge regression (red) compared to the ground truth (blue). The close agreement demonstrates effective regularization.
Table 1: Average MAE (radians) for synthetic spherical trajectory forecasting.
Experiment OLS ($\lambda=0$) Proposed Improvement
ExpGeo 0.035 0.028 20%
ExpCubic 0.043 0.034 21%
ExpSin 0.048 0.037 23%
Average 0.042 0.033 21.4%
3.2 Application: Hurricane Forecasting
Tropical cyclones, also known as hurricanes or typhoons, are among the most powerful natural phenomena with enormous environmental, economic, and human impact. The most common indicator of a hurricane’s intensity is its maximum sustained wind speed, which is used to categorize the storm on the Saffir-Simpson hurricane wind scale. For example, wind speeds $\geq 137$ knots correspond to category 5. High track variability and out-most complexity of hurricanes has led to a large number of works to classify, rationalize and predict them. We remark that many approaches are not intrinsic and use linear approximations. We refer to the overview [OverviewMIHur2020], the summary of recent progress [RecentProgress], and [nhcGuid2022].
In [nava2023hur], we represented hurricane tracks using Bézier splines (compositions of Bézier polynomials), and computed average trajectories, performed tangent principal component analysis, and presented an intensity classification using support vector machines. In the following, we apply our proposed extended ridge regression approach, based on Bézier polynomial representation, to hurricane forecasting.
3.2.1 Dataset
We verify the effectiveness of the proposed framework by applying it to the Atlantic hurricane data from the publicly available  [^1] https://www.nhc.noaa.gov/data/ HURDAT 2 database provided by the U.S. National Oceanic and Atmospheric Administration. The data comprise measurements of latitude, longitude, and wind speed on a 6 hours base. The sample size (number of points that make up a track) varies from 13 to 96. Minimum of observed wind speeds is $15$ knots. Figure 3 illustrates this dataset depicting the 2021 tracks together with their intensities.
3.2.2 Experiments and Discussion
In our experiments, following the verification rules of the U.S. National Hurricane Center (NHC), we computed the average of the MAE (mean absolute error; using spherical distance for the tracks) of the 2021 Atlantic forecasts for comparison. To this end, we represented the tracks and intensities as discrete trajectories in a sphere of radius $3959$ miles (average earth radius) and $\mathbb{R}$, respectively. We initiated each forecast starting with the first sample and choose $n=6$, resulting in average $R^{2}$ values >0.95>0.95 for both fitted tracks and intensities. The high $R^{2}$ values indicate that our polynomials fit well. Higher $n$-values although computationally more intensive, increased the coefficient of determination only incrementally by less then $1\%$. Figure 4 shows two example tracks with their representation in our polynomial regression model.
Figure 4: Two hurricane tracks (blue) with their Bézier polynomial representations (red) using $n=6$ control points. High $R^{2}$ values indicate the polynomial model captures trajectory geometry effectively. These successful fits represent typical cases where our method performs well.
@@ 3.2.2 Experiments and Discussion
We conducted two experiments, Exp1 and Exp2. Due to the deviations of annual trends and relevance of the chronological order, in Exp1, we used the 2020 ones (31 trajectories) as prior and the last 21 for validation to forecast all the 2021 ones (21 trajectories), and in Exp2, the first 16 trajectories from 2021 as prior with its last 5 trajectories to forecast the last 5 trajectories. It turned out that while the intensities are highly correlated, the covariance matrix of the tracks is almost singular (resolved as mentioned before). Higher values of $n$ slightly increased the forecast error of the tracks (overfitting due to small position correlations). For parameter optimization we only employed a simple grid search to decrease the computation time (on average a few seconds for each forecast). Variations in the settings with significantly larger or smaller validation sets slightly increased errors, and, as expected, validation sets chronologically closer to the test sets resulted in forecasts with slightly higher accuracies.
@@
Obviously, one can also model the tracks as images of the Euclidean best-fitting polynomials in geographic coordinates (in the latitude-longitude plane) under the standard parametrization of the sphere. Our computations show that this increases errors ($5\%$ on average) demonstrating the advantage of our intrinsic geometry-aware approach. Figure 5 shows an example with a particularly large error, which is caused by the extreme complexity of dynamics and high variation, particularly near the loop. The increase in error due to the presence of a loop, in addition to meteorological anomalies, is an issue known in other approaches (cf.  [hurRNN2019pred]).
@@
Figure 5: Challenging forecast case: 12 h forecasts (red) via ridge regression for the category 4 and long-lasting 2021 anomalous hurricane Sam (blue). This hurricane has an anomalous highly complex dynamics, and very often and rapidly changes its intensity.
We did not expect our results to match the performance of the NHC’s operational forecasts. Our geometry-aware approach is generic and relies purely on historical trajectory data and statistical-geometric regularization (via Mahalanobis distance), excluding the full complement of physical and atmospheric variables (such as vertical wind shear and sea surface temperature) that are explicitly incorporated into the NHC’s comprehensive suite of dynamical and statistical-dynamical models. Consequently, the NHC results serve as a state-of-the-art performance benchmark rather than a direct, like-for-like comparison. Table 2 summarizes the average 12 h forecast errors of the standard NHC model OCD5 (cf. [nhc2021]) and our two primary experiments regarding the Atlantic basin for the 2021. Our geometric method achieves competitive performance with only a few training trajectories. Deep learning approaches like [hurRNN2019pred] require much more samples and lack interpretability. Our advantage is sample efficiency via geometric priors when training data is scarce (e.g., rare hurricane categories, emerging phenomena). Moreover, our framework currently treats position and intensity independently. A joint representation in $S^{\times}\mathbb{R}$ would capture position-intensity correlations (e.g., rapid intensification over warm water).
Table 2: Average 12-hour forecast MAE for 2021 Atlantic hurricanes.
Experiment Method Intensity (kt) Track (mi)
Exp1 NHC 6.4 48.2
Proposed 6.9 90.7
Exp2 NHC 7.5 67.4
Proposed 6.5 97.2
@@
For the code implementing our approach, which in particular includes Riemannian optimization for the computation of geodesic paths and Bézier polynomials, we used the publicly available Python package morphomatics v4.0 (cf. [Morphomatics]).
4 Conclusions
@@
In this work, we presented an intrinsic natural extension of polynomial ridge regression, respectively, Tikhonov regularization, from Euclidean spaces to arbitrary Riemannian manifolds. This allows for the prediction of manifold-value time series using the Mahalanobis distance to the training data as a prior. We also provided formulae for computing the gradient of the loss function in the approach. Furthermore, due to the use of the de Casteljau algorithm for temporal evaluation of our polynomial model, our approach is computationally fast. We also presented empirical clear error reduction on synthetical spherical data compared to the ordinary least-squared approach. In addition, we provided a discussion of our numerical experiments on the application of the approach to hurricane forecasting and intensity.
Beyond meteorology, our framework applies to: shape trajectories in medical imaging (organ growth, disease progression), rigid body motion in robotics (orientation tracking), covariance evolution in finance (portfolio dynamics), and morphometrics in biology (evolutionary shape changes). By providing explicit gradients and respecting intrinsic geometry, our method advances computational tools for geometric data science, offering sample-efficient alternatives to data-hungry machine learning approaches crucial when training data is scarce.
@@
5 Appendix
5.1 Background: Riemannian Geometry
A Riemannian manifold $M$ is a differentiable manifold endowed with a Riemannian metric $\langle\cdot,\cdot\rangle$, which is a smoothly varying inner product $\langle\cdot,\cdot\rangle_{p}$ on the tangent space $T_{p}M$ for every $p\in M$. This metric induces the Riemannian distance function $dist(p,q)$, defined as the infimum of lengths of curves between $p$ and $q$. If $f:M\to N$ is a smooth map between two manifolds $M$ and $N$, we denote the derivative of $f$ at $p$ in direction $v\in T_{p}M$ by $d_{p}f(v)$. It defines a linear mapping $d_{p}f:T_{p}M\to T_{f(p)}N$. The Levi-Civita connection $\nabla$ applied to vector fields $X$ and $Y$, $\nabla_{X}Y$ is the directional derivative of $Y$ along $X$. The manifold’s generalization of a straight line, is called geodesic (characterized by zero acceleration). A curve $\gamma:I\to M$ is geodesic iff $\nabla_{\gamma^{\prime}}\gamma^{\prime}=0$. For optimization, we rely on tools defined locally within a normal convex neighborhood $U\subset M$, where geodesics are unique and length-minimizing.
Core Geometric Maps
  • Exponential Map ($\exp_{p}$): Maps a tangent vector $v\in T_{p}M$ to the endpoint of the unique geodesic starting at $p$ with initial velocity $v$, i.e., $\exp_{p}(v):=q$.
  • Logarithmic Map ($\log_{p}$): The inverse of $\exp_{p}$, which returns the initial velocity vector $v=\log_{p}(q)\in T_{p}M$ of the geodesic connecting $p$ to $q$.
Calculus and Statistics
The metric structure generalizes necessary calculus and statistical tools for optimization:
  • Gradient: For a smooth function $f:M\to\mathbb{R}$, the gradient at $p$ is defined implicitly by $\d{f}_{p}(v)=\langle\textnormal{grad}_{p}f,v\rangle_{p}$ for all $v\in T_{p}M$. This generalization retains the property that the gradient points in the direction of steepest ascent [riemopt].
  • Fréchet Mean ($\overline{q}$): The sample mean for data $q_{1},\dots,q_{m}\in U$ is generalized as the point that minimizes the sum of squared distances (the Riemannian variance):
    $\overline{q}:=\operatorname{arg\,min}_{p\in U}\sum_{i=1}^{m}dist(p,q_{i})^{2}.$
5.2 Computation of Adjoint Map
Let $x,y\in M$, $v\in T_{x}M$ and $w\in T_{y}M$ and $F:T_{x}M\to T_{y}M$. We recall how to compute $F^{\dagger}$ (the adjoint of $F$).
Using embedding:
Let $A$ denote the matrix representation of $F$ w.r.t. an ONB of the ambient Euclidean space with inner product $\langle\cdot,\cdot\rangle$. Then
$g_{y}(Fv,w)$ $=g_{x}(v,F^{\dagger}w)$
$\langle Fv,w\rangle$ $=\langle v,F^{\dagger}w\rangle$
$(Av)^{T}w$ $=v^{T}(A^{\dagger}w)$
$v^{T}A^{T}w$ $=v^{T}A^{\dagger}w$
Thus, $A^{\dagger}=A^{T}$.
Using local coordinates in a chart:
Let $G_{x}$ and $G_{y}$ denote the (Gram) matrix representations of $g$ in local charts at $x$ and $y$ respectively. Then
$g_{y}(Fv,w)$ $=g_{x}(v,F^{\dagger}w)$
$(Av)^{T}G_{y}w$ $=v^{T}G_{x}(A^{\dagger}w)$
$v^{T}A^{T}G_{y}w$ $=v^{T}G_{x}A^{\dagger}w$
Hence $G_{x}^{-1}A^{T}G_{y}w=A^{\dagger}w$ implying $A^{\dagger}=G_{x}^{-1}A^{T}G_{y}$.
5.3 Algorithmic Forecasting Pipeline
The following summarizes the algorithm for prediction via proposed extended ridge regression. The main objective is to find $b^{*}$ that minimizes the penalized least-squares functional $F_{\lambda}$. Since an explicit solution is generally not available, the minimization is performed iteratively using a Riemannian optimization solver (e.g., gradient descent). We gain the training control points $B$ from training trajectories by fitting polynomials via least-square regression given by $H$. We compute optimal values for the ridge parameter $\lambda$ and the (fine tuning) averaging parameter $\alpha$ using cross-validation and minimizing the residual on a search grid in $0,\infty[\times[0,1]$.
Algorithm 1 Forecasting via Extended Intrinsic Ridge Regression on Manifolds
1:Preprocessing
2:1. Get training and validation sets: Split prior trajectories in training and validation trajectories
3:2. Training: Fit Bézier polynomials to training trajectories $\rightarrow$ control points $B$
4:3. Compute statistics: Calculate Fréchet mean $\mu$ and precision $S$ from $B$
5:4. Optimal $\lambda$ and $\alpha$:
6:  Require search grid $\Gamma=\{(0,0),\cdots,(\lambda_{max},1)\}\subset[0,\infty[\times[0,1]$
7:  Use cross-validation to find optimal values $\lambda^{*}$ and $\alpha^{*}$ by minimizing the residual on $\Gamma$
8:Processing
9:For a new history $(y_{1},\cdots,y_{m})$ predict $y_{m+1}$
10:1. Optimal b via gradient descent: $b^{*}=\underset{b\in N}{\arg\min}F_{\lambda^{*}}(b)$
11:  Require max iterations $K$, learning rate $\eta$, tolerance $\epsilon$
12:  Initialize control points $b^{(0)}$ (e.g., set $b^{(0)}=\mu$)
13:  for $k=0$ to K1K-1
14:  Compute the gradient:
[…]
16:  break
17:  Update the control points using the Riemannian exponential map:
[…]
19:  return $b^{*}=b^{(K)}$
20:2. Result: $\hat{y}_{m+1}=P_{\alpha^{*}}(t_{m+1};b^{*},y_{m})$